OpenStudy (anonymous):

Solve: 4(x − 1) − (x + 2) = 12 5 14 over 3 2 6

OpenStudy (anonymous):

4(x-1)-(x+2)=12 Step 1. Distribute = > 4x-4-x-2=12 Step 2. Simplify => 3x-6=12 Step 3. Combine like terms => 3x=6 Step 4. Solve for x => x=2